# Simultaneous Source Separation, Synchronization, Localization and Mapping for 6G Systems

### v1

@@ Abstract
mpslam is a promising approach for future 6G networks to jointly estimate the positions of transmitters and receivers together with the propagation environment. In cooperative multipath-based simultaneous localization and mapping (MP-SLAM), information collected by multiple mobile terminals is fused to enhance accuracy and robustness. Existing methods, however, typically assume perfectly synchronized base stations and orthogonal transmission sequences, rendering inter-BS interference at the MTs negligible. In this work, we relax these assumptions and address simultaneous source separation, synchronization, and mapping. A relevant example arises in modern 5G systems, where BSs employ muting patterns to mitigate interference, yet localization performance still degrades. We propose a novel BS-dependent data association and synchronization bias model, integrated into a joint Bayesian framework and inferred via the sum-product algorithm on a factor graph. The impact of joint synchronization and source separation is analyzed under various system configurations. Compared with state-of-the-art cooperative MP-SLAM assuming orthogonal and synchronized BSs, our statistical analysis shows no significant performance degradation.
@@ I Introduction
  • Performance analysis demonstrating the impact of joint synchronization and source separation under various system configurations, and comparison with state-of-the-art cooperative MP-SLAM approaches.
@@ III Radio Signal Model and Channel Estimation
Each BS $j$ continuously transmits an radio frequency (RF) pilot signal $s(t)$ of bandwidth $B$ at center frequency $f_{\mathrm{c}}$. Each mobile MT $i$ acts as a receiver, receiving both the LOS component and a set of MPCs arising from reflections off surrounding objects. Since all BSs employ the same pilot sequence, every MT observes an inseparable superposition of signals from all BSs.
@@
${r}_{n}^{(i)}[\ell,h,h^{\prime}]$ $=\sum_{j=1}^{J}\sum_{l=1}^{{N}_{n}^{(j,i)}}{\alpha}_{l,n}^{(j,i)}{e}^{\big(i2\pi\frac{f_{\mathrm{c}}}{c}d^{(j,h)}\cos\big({\vartheta}^{(j,i)}_{l,n}-\psi^{(j,h)}\big)\big)}$
$\hskip 5.69054pt\times{e}^{\big(i2\pi\frac{f_{\mathrm{c}}}{c}d_{n}^{(i,h^{\prime})}\cos\big({\varphi}^{(j,i)}_{l,n}-o_{i,n}-\psi_{n}^{(i,h^{\prime})}\big)\big)}$
@@ IV System Model
At each time $n$, the states of the MTs are given by $\bm{\mathsfbr{x}}_{n}^{(i)}=[\tilde{\bm{\mathsfbr{x}}}^{(i)\hskip 0.85358pt\text{T}}_{n}\hskip 0.85358pt\bm{\mathsfbr{v}}_{n}^{(i)\hskip 0.85358pt\text{T}}~{b}_{\text{mt}\,n}^{(i)}]^{\text{T}}$, where $\tilde{\bm{\mathsfbr{x}}}_{n}^{(i)}$ comprises the position $\bm{\mathsfbr{p}}_{n}^{(i)}$ and the array-orientation $\bm{\mathsfbr{o}}_{n}^{(i)}$ of MT $i$ (for a 2D scenario the dimension of $\mathsfbr{o}_{n}^{(i)}$ is 1 since the orientation is defined by the azimuth angle only). The BS bias is given by ${b}_{\text{bs},n}^{(j)}$ and we also define $\bm{b}_{\text{bs},n}^{(j)}=[{b}_{\text{bs},n}^{(1)}\hskip 2.84526pt\cdots\hskip 2.84526pt{b}_{\text{bs},n}^{(J)}]^{\text{T}}$.
As in [23, 6, 19], we account for the unknown number of VAs by introducing potential virtual anchors (PVAs) $k\in\mathcal{K}_{n}\triangleq\{1,\dots,\mathsfbr{K}_{n}\}$. The number $\mathsfbr{K}_{n}$ of PVAs for all interfering BS $j\in\{1,\dots,J\}$ is the maximum possible number of actual VAs, i.e., all VAs that produced a measurement so far [23, 6, 19] (where $\mathsfbr{K}_{n}$ increases with time). The states of the PVAs are given by $\bm{\mathsfbr{y}}_{k,n}\hskip-0.85358pt=\hskip-0.85358pt\big[\bm{\mathsfbr{p}}^{\text{T}}_{\text{va},k}\;\mathsfbr{r}_{k,n}\big]^{\text{T}}$. The existence/nonexistence of PVAs $k$ is modeled by the existence variable $r_{k,n}\in\{0,1\}$ in the sense that PVAs exists if $r_{k,n}=1$. Formally, its states is considered even if PVAs $k$ is nonexistent, i.e., if $r_{k,n}=0$. The states $\bm{\mathsfbr{p}}_{\text{va},k}$ of nonexistent PVAs are obviously irrelevant. Therefore, all probability density functions defined for PVA states, $f(\bm{y}_{k,n})\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358ptf(\bm{p}_{\text{va},k},r_{k,n})$, are of the form $f(\bm{p}_{\text{va},k},0)\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358ptf_{k,n}f_{\text{d}}(\bm{p}_{\text{va},k})$, where $f_{\text{d}}(\bm{p}_{\text{va},k})$ is an arbitrary “dummy PDF” and $f_{k,n}\!\hskip-0.85358pt\in[0,1]$ is a constant and can be interpreted as the probability of non-existence [23, 6].
@@ IV-A Measurement Model and New PVAs
where the variances, derived from the Fisher information, are $\sigma^{(i)\,2}_{d_{k,n}}=c^{2}/(8\pi^{2}\beta_{\mathrm{bw}}^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n})$, ${\sigma_{\varphi}}^{(k,i)\,2}_{k,n}=1/\!\big(8\pi^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n}D^{2}(\varphi^{(k,i)}_{k,n})\big)$, and ${\sigma_{\vartheta}}^{(k,i)\,2}_{k,n}=1/\!\big(8\pi^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n}D^{2}(\vartheta^{(k,i)}_{l,n})\big)$. Here, $\beta_{\mathrm{bw}}^{2}$ denotes the mean square bandwidth and $D^{2}(\cdot)$ the normalized squared array aperture [24, 21]. Similarly, for the BS-related LHF, we get
@@
Detected VAs, i.e., VAs that generated a measurement for the first time, are modeled by a Poisson process with mean $\mu_{\mathrm{n}}$ and PDF $f_{\mathrm{n}}(\overline{\bm{p}}^{(j,i)}_{\text{va},m}|\bm{x}_{n}^{(i)})$. Accordingly, newly detected VAs are represented by new PVA states $\overline{\bm{\mathsfbr{y}}}^{(j,i)}_{n,m}=[\overline{\bm{\mathsfbr{p}}}^{(j,i)\hskip 0.85358pt\text{T}}_{\text{va},m}\hskip 2.84526pt\overline{\mathsfbr{r}}^{(j,i)}_{m}]^{\text{T}}$, $m\hskip-0.85358pt\hskip-0.85358pt\in\{1,\dots,\mathsfbr{M}^{(i)}_{n}\}$ , $j\in\{1,\dots,J\}$ in our statistical model. Each new PVA state corresponds to a measurement ${\bm{\mathsfbr{z}}_{\text{bs}}}^{(i)}_{m,n}$; $\overline{r}_{m,n}\!=\!1$ implies that measurement ${\bm{\mathsfbr{z}}_{\text{bs}}}^{(i)}_{m,n}$ was generated by a newly detected VA. Since it is unknown which BS (if any) caused a measurement ${\bm{\mathsfbr{z}}_{\text{bs}}}^{(i)}_{m,n}$. Thus, for each index $m$, a number of $J$ new PVAs collected by the stacked vector $\overline{\bm{\mathsfbr{y}}}^{(i)}_{m,n}\triangleq[\overline{\bm{\mathsfbr{y}}}^{(1,i)\hskip 0.85358pt\text{T}}_{m,n}\hskip 0.85358pt\cdots\hskip 0.85358pt\hskip 0.85358pt\hskip 0.85358pt\overline{\bm{\mathsfbr{y}}}^{(J,i)\hskip 0.85358pt\text{T}}_{m,n}]^{\text{T}}$ are generated. The unknown BS index $j$ itself is modeled as a discrete random variable $\overline{\mathsfbr{j}}^{(i)}_{m,n}\in\{1,\dots,J\}$ with uniform prior probability mass function (PMF) $p(\overline{j}^{(i)}_{m,n})\triangleq 1/J$. However, a new PVA can only correspond to one BS resulting in the PDF
@@ IV-B Legacy PVAs and Sequential Update
At time $n$, measurements are incorporated sequentially across MTs $i\hskip-0.85358pt\in\hskip-0.85358pt\{1,\dots,I\}$. Previously detected VAs, i.e., VAs that have been detected either at a previous time $n^{\prime}\!<\hskip-0.85358ptn$ or at the current time $n$ but at a previous MT $i^{\prime}\!<\hskip-0.85358pti$, are therefore represented by legacy PVA states $\underline{\bm{\mathsfbr{y}}}_{k,n}^{(i)}$, with $\underline{\bm{y}}^{(1)}_{k,n}\triangleq\underline{\bm{y}}_{k,n}$.
New PVAs become legacy PVAs when the next measurements—either of the next MT or at the next time instance—are taken into account. In particular, the VA represented by the new VA state $\overline{\bm{\mathsfbr{y}}}^{(j,i^{\prime})}_{m^{\prime},n^{\prime}}$ introduced due to measurement $m^{\prime}$ of MT $i^{\prime}$ corresponding to BS $j$ at $n^{\prime}\hskip-0.85358pt\hskip-0.85358pt\leqslant\hskip-0.85358pt\hskip-0.85358ptn$ is represented by the legacy PVA state $\underline{\bm{\mathsfbr{y}}}^{(i)}_{k,n}$ at time $n$, i.e.,
@@
[…]
fp measurements generated by the CEDA are modeled as a Poisson process with mean $\mu_{\mathrm{fp}}$ and PDF $f_{\mathrm{fp}}(\bm{z}^{(i)}_{m,n})=f_{\mathrm{fp}}({z_{\mathrm{d}}}^{(i)}_{m,n})f_{\mathrm{fp}}({z_{\mathrm{\varphi}}}^{(i)}_{m,n})f_{\mathrm{fp}}({z_{\mathrm{\vartheta}}}^{(i)}_{m,n})$, where the individual FP LHFs are uniformly distributed over $0,d_{\mathrm{max}}]$, $\pi,\pi)$, and $\pi,\pi)$, respectively.
@@ IV-C State Evolution
Legacy PVAs states $\underline{\bm{\mathsfbr{y}}}_{k,n}$ and the MT states $\bm{\mathsfbr{x}}^{(i)}_{n}$ are assumed to evolve independently across time according to state-transition PDFs $f\big(\underline{\bm{y}}_{k,n}\big|\bm{y}_{k,n-1}\big)$, $f(\bm{x}^{(i)}_{n}|\bm{x}^{(i)}_{n-1})$, and $f(b_{\text{bs},n}^{(j)}|b_{\text{bs},n-1}^{(j)})$, respectively. If PVA $k$ exists at time $n\hskip-0.85358pt-\!1$, i.e., $r_{k,n-1}\!=\!1$, it either disappears, i.e., $\overline{r}_{k,n}\!=\hskip-0.85358pt0$, or survives, i.e., $\overline{r}_{k,n}\!=\!1$; in the latter case, it becomes a legacy PVA at time $n$. The probability of survival is denoted by $p_{\mathrm{s}}$. In case of survival, its position remains unchanged, i.e., the state-transition PDF of the VA positions $\underline{\bm{\mathsfbr{p}}}_{\text{va},k}$ is given by $f\big(\underline{\bm{p}}_{\text{va},k}\hskip 0.85358pt\big|\hskip 0.85358pt\bm{p}_{\text{va},k}\big)=\delta\big(\underline{\bm{p}}_{\text{va},k}\hskip-0.85358pt-\hskip 0.85358pt\bm{p}_{\text{va},k}\big)$. Therefore, $f\big(\underline{\bm{p}}_{\text{va},k}\hskip-0.85358pt,\underline{r}_{k,n}\hskip 0.85358pt\big|\hskip 0.85358pt\bm{p}_{\text{va},k},r_{k,n-1}\big)$ for $r_{k,n-1}\hskip-0.85358pt=\hskip-0.85358pt1$ is obtained as
@@
The initial prior PDF at $n\hskip-0.85358pt=\hskip-0.85358pt0$ are given as $f\big(\bm{y}^{(i)}_{k,0}\big)$, $k=\big\{1,\dots,K^{(i)}_{0}\big\}$ and $f(\bm{x}^{(i)}_{0})$, where $\bm{y}^{(i)}_{k,0}$ typically is an empty vector. All (legacy and new) PVA states and all MT states up to time $n$ are denoted as $\bm{\mathsfbr{y}}_{n}\!=\hskip-0.85358pt\big[\bm{\mathsfbr{y}}^{(1)\hskip 0.85358pt\text{T}}_{n}\cdots\hskip 0.85358pt\bm{\mathsfbr{y}}^{(i)\text{T}}_{n}\big]^{\text{T}}$ and $\bm{\mathsfbr{y}}_{0:n}\triangleq\big[\bm{\mathsfbr{y}}^{\text{T}}_{0}\cdots\hskip 0.85358pt\bm{\mathsfbr{y}}^{\text{T}}_{n}\big]^{\text{T}}\!$ and $\bm{\mathsfbr{x}}_{n}=[\bm{\mathsfbr{x}}_{1,n}^{\text{T}}\hskip 0.85358pt\cdots\bm{\mathsfbr{x}}_{I}^{\text{T}}]$ and $\bm{\mathsfbr{x}}_{0:n}=[\bm{\mathsfbr{x}}_{0}^{\text{T}}\hskip 0.85358pt\cdots\bm{\mathsfbr{x}}_{n}^{\text{T}}]$, respectively.

### v2

@@ Abstract
mpslam is a promising approach for future 6G networks to jointly estimate the positions of transmitters and receivers together with the propagation environment. In cooperative multipath-based simultaneous localization and mapping (MP-SLAM), information collected by multiple mobile terminals is fused to enhance accuracy and robustness. Existing methods, however, typically assume perfectly synchronized base stations and orthogonal transmission sequences, rendering inter-BS interference at the MTs negligible. In this work, we relax these assumptions and address simultaneous source separation, synchronization, and mapping. A relevant example arises in modern 5G systems, where BSs employ muting patterns to mitigate interference, yet localization performance still degrades. We propose a novel BS-dependent data association and synchronization bias model, integrated into a joint Bayesian framework and inferred via the sum-product algorithm on a factor graph. The impact of joint synchronization and source separation is analyzed under various system configurations. Compared with state-of-the-art cooperative MP-SLAM assuming orthogonal and synchronized BSs, our statistical analysis shows no significant performance degradation. The proposed BS-dependent data association model constitutes a principled approach for classifying features by arbitrary properties, such as reflection order or feature type (scatterers versus walls).
@@ I Introduction
  • Performance analysis demonstrating the impact of joint synchronization and source separation, and comparison with state-of-the-art cooperative MP-SLAM approaches.
@@ III Radio Signal Model and Channel Estimation
Each BS $j$ continuously transmits an radio frequency (RF) pilot signal $s(t)$ of bandwidth $B$ at center frequency $f_{\mathrm{c}}$. Each MT $i$ acts as a receiver, observing both, the LOS component and MPCs caused by reflections from surrounding objects. Since all BSs employ the same pilot sequence, every MT observes an inseparable superposition of signals from all BSs.
@@
${r}_{n}^{(i)}[\ell,h,h^{\prime}]$ $=\sum_{j=1}^{J}\sum_{l=1}^{{N}_{n}^{(j,i)}}{\alpha}_{l,n}^{(j,i)}{e}^{\big(i2\pi\frac{f_{\mathrm{c}}}{c}d^{(j,h)}\cos({\vartheta}^{(j,i)}_{l,n}-\psi^{(j,h)})\big)}$
$\hskip 5.69054pt\times{e}^{\big(i2\pi\frac{f_{\mathrm{c}}}{c}d_{n}^{(i,h^{\prime})}\cos({\varphi}^{(j,i)}_{l,n}-o_{i,n}-\psi_{n}^{(i,h^{\prime})})\big)}$
@@ IV System Model
At each time $n$, the proposed method infers random variables explicitly denoted in san serif, upright font. Their definitions are in accordance with Sections II and III. The states of the MTs are given by $\bm{\mathsfbr{x}}_{n}^{(i)}=[\tilde{\bm{\mathsfbr{x}}}^{(i)\hskip 0.85358pt\text{T}}_{n}\hskip 0.85358pt\bm{\mathsfbr{v}}_{n}^{(i)\hskip 0.85358pt\text{T}}~\mathsfbr{b}_{\text{mt}\,n}^{(i)}]^{\text{T}}$, with the sub-state $\tilde{\bm{\mathsfbr{x}}}^{(i)\hskip 0.85358pt\text{T}}_{n}=[\bm{\mathsfbr{p}}_{n}^{(i)\hskip 0.85358pt\text{T}}\hskip 0.85358pt\mathsfbr{o}_{n}^{(i)}]^{\text{T}}$. The BS biases $\mathsfbr{b}_{\text{bs},n}^{(j)}$ are collected by $\bm{\mathsfbr{b}}_{\text{bs},n}^{(j)}=[\mathsfbr{b}_{\text{bs},n}^{(1)}\hskip 2.84526pt\cdots\hskip 2.84526pt\mathsfbr{b}_{\text{bs},n}^{(J)}]^{\text{T}}$.
As in [23, 6, 19], we account for the unknown number of VAs by introducing potential virtual anchors (PVAs) $k\in\mathcal{K}_{n}\triangleq\{1,\dots,\mathsfbr{K}_{n}\}$. The number $\mathsfbr{K}_{n}$ of PVAs of all interfering BSs $j\in\{1,\dots,J\}$ is the maximum possible number of actual VAs, i.e., all VAs that produced a measurement so far [23, 6, 19] (where $\mathsfbr{K}_{n}$ increases with time $n$). The states of the PVAs are given by $\bm{\mathsfbr{y}}_{k,n}\hskip-0.85358pt=\hskip-0.85358pt\big[\bm{\mathsfbr{p}}^{\text{T}}_{\text{va},k}\;\mathsfbr{r}_{k,n}\big]^{\text{T}}$. The existence/nonexistence of PVAs $k$ is modeled by the existence variable $r_{k,n}\in\{0,1\}$ in the sense that PVAs exists if $r_{k,n}=1$. Formally, even if PVAs $k$ is nonexistent, i.e., if $r_{k,n}=0$, its state is considered. The states $\bm{\mathsfbr{p}}_{\text{va},k}$ of nonexistent PVAs are obviously irrelevant. Therefore, all probability density functions defined for PVA states, $f(\bm{y}_{k,n})\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358ptf(\bm{p}_{\text{va},k},r_{k,n})$, are of the form $f(\bm{p}_{\text{va},k},0)\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358ptf_{k,n}f_{\text{d}}(\bm{p}_{\text{va},k})$, where $f_{\text{d}}(\bm{p}_{\text{va},k})$ is an arbitrary “dummy PDF” and $f_{k,n}\!\hskip-0.85358pt\in[0,1]$ is a constant and can be interpreted as the probability of non-existence [23, 6].
@@ IV-A Measurement Model and New PVAs
where the means $d^{(k,i)}_{k,n}$, $\varphi^{(k,i)}_{k,n}$ and $\vartheta^{(k,i)}_{k,n}$ are defined in Sec. II, and the variances, derived from the Fisher information, are $\sigma^{(i)\,2}_{d_{k,n}}=c^{2}/(8\pi^{2}\beta_{\mathrm{bw}}^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n})$, ${\sigma_{\varphi}}^{(k,i)\,2}_{k,n}=1/\!\big(8\pi^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n}D^{2}(\varphi^{(k,i)}_{k,n})\big)$, and ${\sigma_{\vartheta}}^{(k,i)\,2}_{k,n}=1/\!\big(8\pi^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n}D^{2}(\vartheta^{(k,i)}_{l,n})\big)$. Here, $\beta_{\mathrm{bw}}^{2}$ denotes the mean square bandwidth and $D^{2}(\cdot)$ the normalized squared array aperture [24, 21]. Similarly, for the BS-related LHF, we get
@@
Detected VAs, i.e., VAs that generated a measurement for the first time, are modeled by a Poisson process with mean $\mu_{\mathrm{n}}$ and PDF $f_{\mathrm{n}}(\overline{\bm{p}}^{(j,i)}_{\text{va},m}|\bm{x}_{n}^{(i)})$. Accordingly, newly detected VAs are represented by new PVA states $\overline{\bm{\mathsfbr{y}}}^{(j,i)}_{n,m}=[\overline{\bm{\mathsfbr{p}}}^{(j,i)\hskip 0.85358pt\text{T}}_{\text{va},m}\hskip 2.84526pt\overline{\mathsfbr{r}}^{(j,i)}_{m}]^{\text{T}}$, $m\hskip-0.85358pt\hskip-0.85358pt\in\{1,\dots,\mathsfbr{M}^{(i)}_{n}\}$ , $j\in\{1,\dots,J\}$ in our statistical model. Each new PVA state corresponds to a measurement ${\bm{\mathsfbr{z}}_{\text{bs}}}^{(i)}_{m,n}$; $\overline{r}_{m,n}\!=\!1$ implies that measurement ${\bm{\mathsfbr{z}}_{\text{bs}}}^{(i)}_{m,n}$ was generated by a newly detected VA. It is also unknown which BS (if any) caused a measurement ${\bm{\mathsfbr{z}}_{\text{bs}}}^{(i)}_{m,n}$. Thus, for each index $m$, a number of $J$ new PVAs collected by the stacked vector $\overline{\bm{\mathsfbr{y}}}^{(i)}_{m,n}\triangleq[\overline{\bm{\mathsfbr{y}}}^{(1,i)\hskip 0.85358pt\text{T}}_{m,n}\hskip 0.85358pt\cdots\hskip 0.85358pt\hskip 0.85358pt\hskip 0.85358pt\overline{\bm{\mathsfbr{y}}}^{(J,i)\hskip 0.85358pt\text{T}}_{m,n}]^{\text{T}}$ are generated. The unknown BS index $j$ itself is modeled as a discrete random variable $\overline{\mathsfbr{j}}^{(i)}_{m,n}\in\{1,\dots,J\}$ with uniform prior probability mass function (PMF) $p(\overline{j}^{(i)}_{m,n})\triangleq 1/J$. However, a new PVA can only correspond to one BS resulting in the PDF
@@
fp measurements generated by the CEDA are modeled as a Poisson process with mean $\mu_{\mathrm{fp}}$ and PDF $f_{\mathrm{fp}}(\bm{z}^{(i)}_{m,n})=f_{\mathrm{fp}}({z_{\mathrm{d}}}^{(i)}_{m,n})f_{\mathrm{fp}}({z_{\mathrm{\varphi}}}^{(i)}_{m,n})f_{\mathrm{fp}}({z_{\mathrm{\vartheta}}}^{(i)}_{m,n})$, where the individual FP LHFs are uniformly distributed over $0,d_{\mathrm{max}}]$, $\pi,\pi)$, and $\pi,\pi)$, respectively. We denote by $\overline{\bm{\mathsfbr{y}}}^{(i)}_{n}\triangleq\big[\hskip 0.85358pt\overline{\bm{\mathsfbr{y}}}^{(i)\hskip 0.85358pt\text{T}}_{1,n}\hskip 0.85358pt\cdots\hskip 0.85358pt\overline{\bm{\mathsfbr{y}}}^{(i)\hskip 0.85358pt\text{T}}_{{\mathsfbr{M}_{\text{bs}}}^{(i)}_{n}\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt,n}\hskip 0.85358pt\big]^{\text{T}}\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt$, the joint vector of all new PVA states.
@@ IV-B Legacy PVAs and Sequential Update
At time $n$, measurements are incorporated sequentially across MTs $i\hskip-0.85358pt\in\hskip-0.85358pt\{1,\dots,I\}$. Previously detected VAs, i.e., VAs that have been detected either at a previous time $n^{\prime}\!<\hskip-0.85358ptn$ or at the current time $n$ but at a previous MT $i^{\prime}\!<\hskip-0.85358pti$, are therefore represented by legacy PVA states $\underline{\bm{\mathsfbr{y}}}_{k,n}^{(i)}$, with $\underline{\bm{\mathsfbr{y}}}^{(1)}_{k,n}\triangleq\underline{\bm{\mathsfbr{y}}}_{k,n}$. Additionally, each legacy PVA implicitly carries an index $j$ that identifies its corresponding BS. It is given by the mapping
$\underline{j}(k)\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt(k\hskip-0.85358pt-\hskip-0.85358pt1)\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\mod\hskip-0.85358pt\hskip-0.85358ptJ+1\,.$ (9)
New PVAs become legacy PVAs when the next measurements—either of the next MT or at the next time instance—are taken into account. In particular, the VA represented by the new VA state $\overline{\bm{\mathsfbr{y}}}^{(j,i^{\prime})}_{m^{\prime},n^{\prime}}$ introduced due to measurement $m^{\prime}$ of MT $i^{\prime}$ corresponding to BS $j$ at $n^{\prime}\hskip-0.85358pt\hskip-0.85358pt\leqslant\hskip-0.85358pt\hskip-0.85358ptn$ is represented by the legacy PVA state $\underline{\bm{\mathsfbr{y}}}^{(i)}_{k,n}=[\underline{\bm{\mathsfbr{p}}}^{(i)\hskip 0.85358pt\text{T}}_{\text{va},m}\hskip 2.84526pt\underline{\mathsfbr{r}}^{(i)}_{m}]^{\text{T}}$ at time $n$, i.e.,
@@ IV-C State Evolution
Legacy PVAs states $\underline{\bm{\mathsfbr{y}}}_{k,n}$ and the MT states $\bm{\mathsfbr{x}}^{(i)}_{n}$ are assumed to evolve independently across time according to state-transition PDFs $f(\bm{x}^{(i)}_{n}|\bm{x}^{(i)}_{n-1})$, $f(b_{\text{bs},n}^{(j)}|b_{\text{bs},n-1}^{(j)})$, and $f\big(\underline{\bm{y}}_{k,n}\big|\bm{y}_{k,n-1}\big)$, respectively. In particular, if PVA $k$ exists at time $n\hskip-0.85358pt-\!1$, i.e., $r_{k,n-1}\!=\!1$, it either disappears, i.e., $\overline{r}_{k,n}\!=\hskip-0.85358pt0$, or survives, i.e., $\overline{r}_{k,n}\!=\!1$; in the latter case, it becomes a legacy PVA at time $n$. The probability of survival is denoted by $p_{\mathrm{s}}$. In case of survival, its position remains unchanged, i.e., the state-transition PDF of the VA positions $\underline{\bm{\mathsfbr{p}}}_{\text{va},k}$ is given by $f\big(\underline{\bm{p}}_{\text{va},k}\hskip 0.85358pt\big|\hskip 0.85358pt\bm{p}_{\text{va},k}\big)=\delta\big(\underline{\bm{p}}_{\text{va},k}\hskip-0.85358pt-\hskip 0.85358pt\bm{p}_{\text{va},k}\big)$. Therefore, $f\big(\underline{\bm{p}}_{\text{va},k}\hskip-0.85358pt,\underline{r}_{k,n}\hskip 0.85358pt\big|\hskip 0.85358pt\bm{p}_{\text{va},k},r_{k,n-1}\big)$ for $r_{k,n-1}\hskip-0.85358pt=\hskip-0.85358pt1$ is obtained as
@@
The initial prior PDF at $n\hskip-0.85358pt=\hskip-0.85358pt0$ are given as $f\big(\bm{y}^{(i)}_{k,0}\big)$, $k=\big\{1,\dots,K^{(i)}_{0}\big\}$ and $f(\bm{x}^{(i)}_{0})$, where $\bm{\mathsfbr{y}}^{(i)}_{k,0}$ typically is an empty vector. All (legacy and new) PVA states and all MT states up to time $n$ are denoted as $\bm{\mathsfbr{y}}_{n}\!=\hskip-0.85358pt\big[\bm{\mathsfbr{y}}^{(1)\hskip 0.85358pt\text{T}}_{n}\cdots\hskip 0.85358pt\bm{\mathsfbr{y}}^{(i)\text{T}}_{n}\big]^{\text{T}}$ and $\bm{\mathsfbr{y}}_{0:n}\triangleq\big[\bm{\mathsfbr{y}}^{\text{T}}_{0}\cdots\hskip 0.85358pt\bm{\mathsfbr{y}}^{\text{T}}_{n}\big]^{\text{T}}\!$ and $\bm{\mathsfbr{x}}_{n}=[\bm{\mathsfbr{x}}_{1,n}^{\text{T}}\hskip 0.85358pt\cdots\bm{\mathsfbr{x}}_{I}^{\text{T}}]$ and $\bm{\mathsfbr{x}}_{0:n}=[\bm{\mathsfbr{x}}_{0}^{\text{T}}\hskip 0.85358pt\cdots\bm{\mathsfbr{x}}_{n}^{\text{T}}]$, respectively. To model the sequential update of the BS biases, we define the factors $f^{(i)}(b_{\text{bs},n}^{(j,i)}|b_{\text{bs},n}^{(j,i-1)})$ for $i\geqslant 2$ as $f^{(i)}(b_{\text{bs},n}^{(j,i)}|b_{\text{bs},n}^{(j,i-1)})=\delta(b_{\text{bs},n}^{(j,i)}-b_{\text{bs},n}^{(j,i-1)})$ with $b_{\text{bs},n}^{(j)}\triangleq b_{\text{bs},n}^{(j,I)}$. For $i=1$, we use $f^{(1)}(b_{\text{bs},n}^{(j,1)}|b_{\text{bs},n}^{(j,0)})$ with $b_{\text{bs},n}^{(j,0)}\triangleq b_{\text{bs},n-1}^{(j)}$.